OpenStudy (anonymous):

Find the perimeter of the rectangle in simplest form. 4 + √6 and 2√6

Parth (parthkohli):

What is the width, and what's the length?

OpenStudy (anonymous):

2 radical 6 is the length and the other the width

Parth (parthkohli):

4 + radical 6 is the width?

OpenStudy (anonymous):

Yup

Parth (parthkohli):

Okay, so the perimeter of a rectangle is given by \(2(l + w)\).

Parth (parthkohli):

So, \(\Large \color{Black}{\Rightarrow 2(4 + \sqrt{6} + 2\sqrt{6}) }\) \(\Large \color{Black}{\Rightarrow 2(4 + 3\sqrt{6}) }\) \(\Large \color{Black}{\Rightarrow 2(4) + 2(3 \sqrt{6}) }\) \(\Large \color{Black}{\Rightarrow 8 + 6\sqrt{6} }\)